What is the pH of a buffer prepared by adding 6.0 mol of acetic acid and 1 mol of NaOH to water to make a 1.0-L solution? Ka for acetic acid is 1.75*10-5 and its pKa is 4.76.
Answer #1

pH= pka + log [acetate ]/[acetic acid]

acetic acid moles = 6

NaOOH moles added = 1

acetic acid reacts with NaOH to form sodium acetate

henceacetic acid after reaction = 6-1 = 5 ,

acetate moles = NaoH moels reacted = 1

Now [acetic acdi] = moles/vol = 5/1 = 5 , [acetate] = 1/1 = 1 ,

now pH = 4.76 + log ( 1/5)

     = 4.06